OpenStudy (aravindg):

(m+1)x^2+2mx+3=0 sum of roots 1 m=?

OpenStudy (anonymous):

For myininaya http://openstudy.com/groups/mathematics/updates/4e41d1280b8b1fdcd4b52b71

myininaya (myininaya):

\[b=\frac{-2m \pm \sqrt{(2m)^2-4(m+1)(3)}}{2(m+1)}=\frac{-2m \pm \sqrt{4m^2-12m-12}}{2(m+1)}\] \[=-\frac{-2m \pm \sqrt{4(m^2-3m-3)}}{2(m+1)}=\frac{-2m \pm 2\sqrt{m^2-3m-3}}{2(m+1)}\] \[=\frac{-m \pm \sqrt{m^2-3m-3}}{m+1}\] \[\frac{-m +\sqrt{m^2-3m-3}}{m+1}+\frac{-m - \sqrt{m^2-3m-3}}{m+1}=1\]

myininaya (myininaya):

\[\frac{-2m}{m+1}=1\]

myininaya (myininaya):

\[-2m=m+1, m \neq -1\]

myininaya (myininaya):

\[-2m-m=1 => -3m=1 => m=\frac{-1}{3}\]

OpenStudy (anonymous):

Oh dear, no alpha + beta = -b/a (standard formula)

OpenStudy (aravindg):

so wat is answer?

OpenStudy (anonymous):

Am I misunderstanding the question?

myininaya (myininaya):

sum of roots is 1 right?

OpenStudy (aravindg):

yes

myininaya (myininaya):

ok oh and also estudier -2m/(m+1) is -b/a

myininaya (myininaya):

and this is suppose to be one according to what he has

OpenStudy (aravindg):

answer is -1/3?

myininaya (myininaya):

if i made no mistake

myininaya (myininaya):

but i'm known for making alot of mistakes maybe you can trust me and maybe you can't who knows

myininaya (myininaya):

but it looks good to me

OpenStudy (aravindg):

i truust u

myininaya (myininaya):

lol

OpenStudy (anonymous):

U can go do my question now:-)

OpenStudy (aravindg):

lol

myininaya (myininaya):

ok i'm gonna get on my laptop and move to the other room i'm about to watch white collar

OpenStudy (aravindg):

u hav a fan